OpenStudy (anonymous):

Find the slope of the line. 2y-6x=5 Mental Math

OpenStudy (marigirl):

why dont u make y the subject of the equation. making it as y=mx+c form

OpenStudy (anonymous):

2y-6x=5 2y=6x+5 y=3x+5/2 slope equals 3